OpenStudy (anonymous):

How would I graph y=6|x-7|?

OpenStudy (anonymous):

hint I x I = -(x) , x<0 =x , x>0

OpenStudy (anonymous):

so do the same for I x-7 I I x-7 I = -(x-7) , x<7 = (x-7), x>7

OpenStudy (anonymous):

What about the 6?

OpenStudy (anonymous):

now u have two equation...plot them y=-6(x-7) , x<7 y=6 (x-7) , x>7

OpenStudy (anonymous):

Oh, okay. Thankyou! It really helped

OpenStudy (anonymous):

:)